OpenStudy (anonymous):

Which ordered pair in the form (a, b) is a solution of this equation? 4a – 5b = 24 A. (–2, –5) B. (0, 6) C. (–5, –2) D. (6, 0)

OpenStudy (vrajrpatel):

B

OpenStudy (kayders1997):

No, when doing this I would just plug in your a value from the coordinate and you b value from the coordinate and see if they add up to the total

OpenStudy (kayders1997):

Should I just you how to do what I am explaining with the option B?

OpenStudy (anonymous):

yeah

OpenStudy (kayders1997):

okay the coordinate is (0,6) and it says that it is in the form (a,b) so we know the 0 is a and the 6 is b let's put that into our equation 4a-5b=24 so 4(0)-5(6)=24 just plugging in a and b so 0-30=24 -30 does not equal 24

OpenStudy (kayders1997):

@swagmadness251